# Autonomous Stabilization of Floquet States Using Static Dissipation

### v2

@@ V Dissipative Stabilization of a Topological Photon Pump
The boost cavity—with frequency $\Delta_{\mathrm{b}}$, annihilation operator $a_{\mathrm{b}}$, and coupling $g_{\mathrm{b}}$ to the qubit—holds the non-classical state which is to be boosted, and ideally would be lossless. The stabilizing cavity—with label $\mathrm{s}$ to differentiate it from the boost cavity—stabilizes the resource qubit. It experiences photon loss with a power decay rate $\kappa_{\mathrm{s}}$ (which we model through Lindblad evolution with a jump operator $a_{\mathrm{s}}$). We choose a qubit drive
@@
Cavity state boosting (and topological photon pumping more broadly Nathan et al. (2019)) requires the boost cavity energy and drive frequencies to be much smaller than the qubit energy and interaction: $\Delta_{\mathrm{b}},\omega_{\mathrm{mod}}\ll B_{0},g_{\mathrm{b}}\sqrt{n_{\mathrm{b}}}$. Meanwhile, stabilization requires a resonance condition between the qubit quasienergy (which can be replaced by the instantaneous energy near the adiabatic limit) and the stabilizing cavity: $|\Delta_{\mathrm{s}}-B_{0}|\lesssim g_{\mathrm{s}}$. In order for the boost cavity to not couple strongly to the stabilizing cavity, and inherit its photon loss, we also demand that they are far detuned: $|\Delta_{\mathrm{s}}-\Delta_{\mathrm{b}}|\gg g_{\mathrm{b}},g_{\mathrm{s}}$.
@@
The introduction of the auxiliary stabilizing cavity suppresses nonadiabatic excitation of the qubit along the curved part of the semicircle protocol through precisely the same mechanism as Sec. III. Without this, the qubit has a higher probability to flip to the |$\ket{g}$ state by the beginning of the vertical segment, and then absorb photons from the cavity by adiabatically following the opposite branch of the avoided crossing, transitioning from |$\ket{g,n}$ to |$\ket{e,n-1}$.
We numerically verify the quantitative improvement of introducing coupling to the auxiliary lossy cavity in Figs. 5(b,c) using a Lindblad master equation simulation. The stabilized boosting protocol produces an average pumping rate which is closer to the expected value of one photon per period [Eq. (13)] and a boosted Fock state with a smaller variance, despite the introduction of dissipation. This is shown quantitatively in Appendix D, where we also discuss details of our numerical simulations.
Figure 5: Stabilized cavity state boosting. (a) The $\vec{B}$ field for the boosting protocol traces out a semicircle of radius $B_{0}$ (black) on the $B_{x}-B_{z}$ plane with period $T_{\mathrm{mod}}$. On the blue dashed circle, the auxiliary cavity s is resonant with the qubit, while on the red dashed circle, the boost cavity b is resonant with the qubit. (b, c) Plots of the Fock state occupation of the boost cavity ($P(n_{\mathrm{b}})=\bra{n_{\mathrm{b}}}\rho_{\mathrm{b}}(t)\ket{n_{\mathrm{b}}}$ where $\rho_{\mathrm{b}}(t)$ is the reduced density matrix for the boost cavity) without (b) and with (c) the auxiliary cavity as a function of protocol duration. In (b), nonadiabatic excitations of the qubit produce significant population in the cavity at small photon number, indicated by a dark color in the figure. As a result, the expected photon number in the cavity state (blue dashed) is significantly below the quantized value (red dashed). (c) With the auxiliary cavity, nonadiabatic excitations of the qubit are suppressed, resulting in $\langle n_{b}\rangle$ becoming closer to the quantized value and cavity states that are closer to the desired high photon number Fock state at the special rephasing times $t^{*}$. Parameters: $B_{0}/g_{\mathrm{b}}=20$, $\omega_{\mathrm{mod}}/g_{\mathrm{b}}=1.5$, $\Delta_{\mathrm{b}}/\omega_{\mathrm{mod}}=(1+\sqrt{5})/2$, $\Delta_{\mathrm{s}}/B_{0}=1$, initial state |$\ket{g,n_{\mathrm{b}}=10,n_{\mathrm{s}}=0}$; (b) $g_{\mathrm{s}}/g_{\mathrm{b}}=0$; (c) $g_{\mathrm{s}}/g_{\mathrm{b}}=1$, photon loss from the lossy cavity with rate $\kappa_{\mathrm{s}}/g_{\mathrm{b}}=1$ is modeled with a jump operator $a_{\mathrm{s}}$ in a Lindblad equation, other dissipative processes are neglected.
@@ VI Conclusion and Outlook
However, there are two aspects of our results that suggest that the same scheme can combat these other kinds of heating. First, our experimental implementation is carried out on a transmon qubit, which is not a true two-level system. In fact, the transmon is notorious for being only loosely describable as a two-level system due to its relatively low anharmonicity of $100-250$ MHz (or roughly $2\%-5\%$). The data in Fig. 2 shows persistent oscillations for $100\ \mu$s and over $100$ drive periods without signs of uncompensated heating due to the higher levels of the transmon, despite applied drive strengths which are roughly one third of the anharmonicity. Second, the stabilized topological photon pump described in Sec. V uses a second lossy cavity to stabilize a Floquet system consisting of both a qubit and a cavity. The presence of autonomous stabilization in both of these cases despite the larger Hilbert space is a strong indicator that it is possible to suppress Floquet heating in general using our simple scheme, rather than the contributions of excited-state lifetime alone. There is also evidence that the scheme can combat heating due to many-body interactions, and can cool into desired ground states of effective Floquet Hamiltonians Petiziol and Eckardt (2022); Schaefer et al. . However, a full characterization of the capability of autonomous stabilization with a static auxiliary remains an outstanding question.
Optimal stabilization, with the fastest possible stabilization time and highest possible steady-state purity, relies on exact resonances in the augmented spin-cavity system. However, significant stabilization can be achieved for a much wider range of parameters. Supplementary experimental data, shown in Appendix F, shows that significant stabilization is observed for all configurations in the range $|\delta|\lesssim g$. As the transmon decay rate is only a fourth of the photon decay rate, using an even lossier cavity (up to $\kappa\approx g$) would result in stabilization for an even wider range of parameters. Additionally, we find that elliptical drive fields, which are only resonant during part of the drive cycle, also exhibit stabilization. The combination of these two experimental observations is a strong indicator that a single lossy auxiliary degree of freedom can be used to counteract heating into all excited states within a range of roughly $g$ of the resonance, and hence systematically remove population from entire targeted Floquet bands. Alternatively, an auxiliary consisting of a band-pass filter, rather than a single cavity, could be used to enhance loss for a large number of Floquet states Wang and Safavi-Naeini (2023). Implementations along these lines would give the ability to, for example, cool directly into the ground band of a Floquet engineered lattice Seetharam et al. (2015), rather than relying on adiabatic band mapping, which is generally only possible in synthetic lattice systems.
@@ Acknowledgements.
We thank Maya Amouzegar for assistance with device fabrication, Benjamin Cochran for contributions to device modeling, IBK Adisa for assistance with early measurements; and Vedika Khemani, Chris Laumann, and Paul Schindler for helpful comments. During completion of this manuscript, we became aware of an unrelated proposal Lin et al. (2024) in which the ability to achieve resonance in a Jaynes-Cummings model using strong Rabi drives was harnessed for a topological pump. This work was supported by ARL (Grants no. W911NF-19-2-0181 and W911NF-17-S-0003), the University of Maryland, NSF (QLCI grant OMA-2120757), and AFOSR (Grants No. FA9550-20-1-0235 and FA9550-24-1-0121). MR received support from the National Science Foundation (PFC at JQI Grant No. PHY-1430094), the Laboratory for Physical Sciences graduate fellowship, and ARCS. QY received support from AFOSR Grant No. FA9550-21-1-0129. DL was supported by the Laboratory for Physical Sciences, a Stanford Q-FARM Bloch Postdoctoral Fellowship, and the Packard Foundation through a Packard Fellowship in Science and Engineering.
@@ Appendix B Synthesizing a Rotating Field
We highlight this difference by explicitly comparing the two rotating-frame Hamiltonians obtained for a qubit coupled to a cavity. The general laboratory-frame Hamiltonian which incorporates both frequency-modulation of the qubit and the drive is:
$H=\omega_{m}a^{\dagger}a+\omega_{q}(t)\frac{\sigma_{z}}{2}+E_{0}(t)\cos[\omega_{d}(t)\times t](\sigma^{-}+\sigma^{+})+g(a^{\dagger}\sigma^{-}+a\sigma^{+}).$ (21)
If the drive frequency is held fixed, and only the qubit frequency is varied, the effective rotating frame Hamiltonian is simplest in the non-time-dependent rotating frame of the drive:
$H_{\mathrm{eff}}=\Delta_{dc}a^{\dagger}a+\frac{\Delta_{dq}(t)}{2}\sigma_{z}+\frac{\Omega(t)}{2}\sigma_{x}+g(a^{\dagger}\sigma^{-}+a\sigma^{+})$ (22)
where $\Delta_{dc}=\omega_{m}-\omega_{d}$ is the cavity-drive detuning and $\Delta_{dq}(t)=\omega_{q}(t)-\omega_{d}$ is the qubit-drive detuning. In this configuration, the qubit-cavity detuning picks up a time varying component: $\Delta_{qc}=\omega_{m}-\omega_{q}=\Delta_{dc}-\Delta_{dq}(t)$. However, if the qubit frequency is held fixed and the drive frequency is varied, then the rotating-frame Hamiltonian is simplest in the rotating frame of the qubit instead:
$H_{\mathrm{eff}}=\Delta_{qc}a^{\dagger}a+\frac{\Delta_{dq}(t)}{2}\sigma_{z}+\frac{\Omega(t)}{2}\sigma_{x}+g(a^{\dagger}\sigma^{-}+a\sigma^{+})$ (23)
where $\Delta_{qc}=\omega_{m}-\omega_{q}$ is now explicitly time independent, and $\Delta_{dq}(t)=\omega_{q}-\omega_{d}(t)$.
@@ Appendix D Numerical simulations of the augmented spin-cavity and boosting setups
with $L_{1}=\sqrt{\kappa}a$ (the jump operator modelling photon loss), $L_{2}=\sqrt{\Gamma}\sigma^{-}$ (the jump operator modelling qubit loss, or $T_{1}$ relaxation), and $\{A,B\}=AB+BA$ being the anticommutator. In Sec. 5 we use a similar Lindblad framework in a Hilbert space with two cavities, but with $H_{B}^{+}$ replaced by $H_{\mathrm{boost}}$ and a single jump operator $L_{1}=\sqrt{\kappa_{\mathrm{s}}}a_{\mathrm{s}}$. We solve for $\rho(t)$ given some initial conditions by direct integration in a truncated cavity Hilbert space (we have checked that the state remains far from the truncation throughout the evolution), using the master equation methods of QuTiP Johansson et al. (2013). The data displayed in Figs. 5, 9, and 11 is obtained in this way.
@@ Appendix E Onset of stabilization in fast ramp-out measurements
Figure 10: Onset of stabilization. (a) The probability of being in state |$\ket{e}$ as a function of time for different value of minimum detuning to cavity resonance $\delta/2\pi=(\Delta-B_{0})/2\pi$ measured using the fast ramp out protocol. The orange (blue) curves are obtained by starting with the qubit aligned (anti-aligned) with the field, which we denote as |$\ket{\phi_{+}}_{0}$ and |$\ket{\phi_{-}}_{0}$, respectively. (a.i) Far from resonance ($\delta\gg g$), both state initializations damp out to a mixed state while maintaining a $\pi$-phase shift. (a.ii-a.iii) As the detuning $\delta$ decreases, the state initially prepared in |$\ket{\phi_{+}}$ damps out. Its signal then grows to a non-zero oscillatory amplitude in phase with that from the |$\ket{\phi_{-}}$ state. (a.iv) On resonance, both states are quickly stabilized to |$\ket{\phi_{-}(t)}$. (b) Corresponding data acquired using the slow ramp out protocol, effectively showing the envelope of the oscillatory data in panels (a.i-a.iv). Curves are labeled by the minimum detuning to cavity resonance $\delta/2\pi$ in MHz. Exponential fits to the data are shown in dashed (dotted) lines for preparation in |$\ket{\phi_{-}}$ (|$\ket{\phi_{+}}$). Insets on the right of (a) show a schematic representation of the resonance condition in the $B_{x}-B_{z}$ plane. The solid black line denotes the magnetic field strength and the dashed line denotes the mean qubit-cavity detuning (the colored bands show the ±$\pm g$ width from the cavity-qubit coupling).
In addition to the slow ramp-out data presented in Fig. 3, it is also possible to observe the onset of stabilization in the bare qubit |$\ket{g}$, |$\ket{e}$ basis using the fast ramp out protocol. This protocol gives direct access to the time dependence of the instantaneous eigenstates. States prepared in |$\ket{\phi_{-}}$ (using an initial $\pi$-pulse before the protocol starts) evolve with a $\pi$-phase shift relative to states initialized in |$\ket{\phi_{+}}$ as shown in the blue and orange traces in panels (a.i)-(a.ii) in Fig. 10 respectively. As $\delta=\Delta-B_{0}$ approaches zero, states initially prepared in |$\ket{\phi_{+}}$ damp out before being stabilized in the |$\ket{\phi_{-}}$ state, oscillating at high contrast indefinitely. The instantaneous eigenstate data [shown again in panel (b) for convenience], represents the |$\ket{\phi_{-}}$ population for the oscillatory data presented in panels (a.i)-(a.iv) with 0.5 indicating a perfectly mixed state (i.e. zero oscillation contrast).

### v3

@@ V Dissipative Stabilization of a Topological Photon Pump
The boost cavity—with frequency $\Delta_{\mathrm{b}}$, annihilation operator $a_{\mathrm{b}}$, and coupling $g_{\mathrm{b}}$ to the qubit—holds the non-classical state which is to be boosted, and ideally would be lossless. The stabilizing cavity—with label $\mathrm{s}$ to differentiate it from the boost cavity—stabilizes the resource qubit. It experiences photon loss with a power decay rate $\kappa_{\mathrm{s}}$ (which we model through Lindblad evolution with a jump operator $a_{\mathrm{s}}$).
We choose a qubit drive
@@
Figure 5: Stabilized cavity state boosting. (a) The $\vec{B}$ field for the boosting protocol traces out a semicircle of radius $B_{0}$ (black) on the $B_{x}-B_{z}$ plane with period $T_{\mathrm{mod}}$. On the blue dashed circle, the auxiliary cavity s is resonant with the qubit, while on the red dashed circle, the boost cavity b is resonant with the qubit. (b, c) Plots of the Fock state occupation of the boost cavity ($P(n_{\mathrm{b}})=\bra{n_{\mathrm{b}}}\rho_{\mathrm{b}}(t)\ket{n_{\mathrm{b}}}$ where $\rho_{\mathrm{b}}(t)$ is the reduced density matrix for the boost cavity) without (b) and with (c) the auxiliary cavity as a function of protocol duration. In (b), nonadiabatic excitations of the qubit produce significant population in the cavity at small photon number, indicated by a dark color in the figure. As a result, the expected photon number in the cavity state (blue dashed) is significantly below the ideal, quantized, value [red dashed, Eq. (13)]. (c) With the auxiliary cavity, nonadiabatic excitations of the qubit are suppressed, resulting in $\langle n_{b}\rangle$ becoming closer to the quantized value and cavity states that are closer to the desired high photon number Fock state at the special rephasing times $t^{*}$. Parameters: $B_{0}/g_{\mathrm{b}}=20$, $\omega_{\mathrm{mod}}/g_{\mathrm{b}}=1.5$, $\Delta_{\mathrm{b}}/\omega_{\mathrm{mod}}=(1+\sqrt{5})/2$, $\Delta_{\mathrm{s}}/B_{0}=1$, initial state |$\ket{g,n_{\mathrm{b}}=10,n_{\mathrm{s}}=0}$; (b) $g_{\mathrm{s}}/g_{\mathrm{b}}=0$; (c) $g_{\mathrm{s}}/g_{\mathrm{b}}=1$, photon loss from the lossy cavity with rate $\kappa_{\mathrm{s}}/g_{\mathrm{b}}=1$ is modeled with a jump operator $a_{\mathrm{s}}$ in a Lindblad equation, other dissipative processes are neglected.
Cavity state boosting (and topological photon pumping more broadly) requires the boost cavity energy and drive frequencies to be much smaller than the qubit energy and interaction: $\Delta_{\mathrm{b}},\omega_{\mathrm{mod}}\ll B_{0},g_{\mathrm{b}}\sqrt{n_{\mathrm{b}}}$ Nathan et al. (2019). Meanwhile, stabilization requires a resonance condition between the qubit quasienergy (which can be replaced by the instantaneous energy near the adiabatic limit) and the stabilizing cavity: $|\Delta_{\mathrm{s}}-B_{0}|\lesssim g_{\mathrm{s}}$.
In order for the boost cavity to not couple strongly to the stabilizing cavity, and inherit its photon loss, we also demand that they are far detuned: $|\Delta_{\mathrm{s}}-\Delta_{\mathrm{b}}|\gg g_{\mathrm{b}},g_{\mathrm{s}}$.
@@
The introduction of the auxiliary stabilizing cavity suppresses nonadiabatic excitation of the qubit along the curved part of the semicircle protocol through precisely the same mechanism as Sec. III. Without this, the qubit has a higher probability to flip to the |$\ket{g}$ state by the beginning of the vertical segment, and then absorb photons from the cavity by adiabatically following the opposite branch of the avoided crossing, transitioning from |$\ket{g,n}$ to |$\ket{e,n-1}$. This nonadiabatic process results in a cavity state which: (i) has a lower mean photon number than predicted by the ideal, quantized, rate of energy pumping in Eq. (13); and (ii) is no longer a perfect Fock state, instead having some nonzero variance in its photon number distribution.
We numerically verify the quantitative improvement of boosting due to introducing coupling to the auxiliary lossy cavity in Figs. 5(b,c) using a Lindblad master equation simulation. The stabilized boosting protocol produces (i) an average pumping rate which is closer to the quantized expected value of one photon per period [Eq. (13)]; and (ii) a boosted Fock state with a smaller photon number variance, despite the introduction of dissipation. This is shown quantitatively in Appendix D, where we also discuss details of our numerical simulations.
@@ VI Conclusion and Outlook
However, there are two aspects of our results that suggest that the same scheme can combat these other kinds of heating. First, our experimental implementation is carried out on a transmon qubit, which is not a true two-level system. In fact, the transmon is notorious for being only loosely describable as a two-level system due to its relatively low anharmonicity of $100-250$ MHz (or roughly $2\%-5\%$). The data in Fig. 3 shows persistent oscillations for $100\ \mu$s and over $100$ drive periods without signs of uncompensated heating due to the higher levels of the transmon, despite applied drive strengths which are roughly one third of the anharmonicity. Second, the stabilized topological photon pump described in Sec. V uses a second lossy cavity to stabilize a Floquet system consisting of both a qubit and a cavity. The presence of autonomous stabilization in both of these cases despite the larger Hilbert space is a strong indicator that it is possible to suppress Floquet heating in general using our simple scheme, rather than the contributions of excited-state lifetime alone. However, a full characterization of the capability of autonomous stabilization with a static auxiliary remains an outstanding question.
A major advantage of dissipative stabilization methods, such as those we present here, is that they are robust to slight imperfections and variations of parameters. While optimal stabilization—with the fastest possible stabilization time and highest possible steady-state purity—relies on exact resonances in the augmented spin-cavity system, significant stabilization can be achieved for a much wider range of parameters. This includes both varying parameters within our studied driving protocol, but also altering the protocol completely, or changing experimental systems entirely. Indeed, supplementary experimental data (Appendix F) shows that significant stabilization is observed for all drive parameters in the range $|\delta|\lesssim g$. As the transmon decay rate in our device is only a fourth of the photon decay rate, using an even lossier cavity (up to $\kappa\approx g$) would result in stabilization for an even wider range of detunings. Varying our drive protocol, we also find that elliptical drive fields, which are only instantaneously near-resonant during part of the drive cycle, also exhibit stabilization.
The robustness of dissipative stabilization opens the door to potential applications in settings far beyond the experiment performed here. For instance, a single sufficiently lossy auxiliary degree of freedom could be used to counteract heating into all excited states within a range of roughly $g$ of the resonance, and hence systematically remove population from targeted Floquet bands. If either the Floquet band gap or the cavity frequency can be varied in time, then the entire Floquet band could be emptied of excitations. Using similar time-dependent schemes, entangled states can be prepared, simplifying current schemes that target multiple resonances at the same time Shankar et al. (2013); Brown et al. (2022).
Alternatively, an auxiliary consisting of a band-pass filter, rather than a single cavity, could be used to enhance loss for a large number of Floquet states Wang and Safavi-Naeini (2023). Implementations along these lines would give the ability to, for example, cool directly into the ground state of a Floquet engineered Hamiltonian Seetharam et al. (2015), rather than relying on adiabatic state preparation, which is slow and fails near phase transitions. We note that our natively-Floquet dissipative scheme is well-adapted to emerging qubit designs which utilize periodic driving explicitly Wang and Safavi-Naeini (2023).
When applied to a many-body periodically driven system in the high-frequency regime, the scheme discussed in this article can cool the system to close to the ground state of a gapped effective Hamiltonian Petiziol and Eckardt (2022); Schäefer et al. . The effective Hamiltonian controls stroboscopic dynamics and can be derived in a high-frequency expansion Bukov et al. (2015). In the absence of the dissipative element, the time evolution of the system is captured by a Gibbs state with a slowly increasing temperature as a function of time. The dissipative element removed entropy from this Gibbs state and combats the slow heating for optimal choices of $g,\kappa$. When the effective Hamiltonian is gapped, our upcoming work Schäefer et al. shows that low temperatures smaller than the gap of the effective Hamiltonian can be reached in the steady state.
@@ Acknowledgements.
We thank Maya Amouzegar for assistance with device fabrication, Benjamin Cochran for contributions to device modeling, IBK Adisa for assistance with early measurements; and Vedika Khemani, Chris Laumann, and Paul Schindler for helpful comments. During completion of this manuscript, we became aware of an unrelated proposal Lin et al. (2024) in which the ability to achieve resonance in a Jaynes-Cummings model using strong Rabi drives was harnessed for a topological pump. This work was supported by ARL (Grants no. W911NF-19-2-0181 and W911NF-17-S-0003), the University of Maryland, and AFOSR (Grants No. FA9550-20-1-0235 and FA9550-24-1-0121). MR received support from the National Science Foundation (PFC at JQI Grant No. PHY-1430094), the Laboratory for Physicsal Sciences graduate fellowship, and ARCS. QY received support from AFOSR Grant No. FA9550-21-1-0129. DL was supported by the Laboratory for Physical Sciences, a Stanford Q-FARM Bloch Postdoctoral Fellowship, and the Packard Foundation through a Packard Fellowship in Science and Engineering (PI: Vedika Khemani). AC thanks the Max Planck Institute for the Physics of Complex Systems for its hospitality
@@ Appendix B Synthesizing a Rotating Field
We highlight this difference by explicitly constructing the rotating-frame Hamiltonian obtained for a qubit coupled to a cavity. The general laboratory-frame Hamiltonian which incorporates both frequency-modulation of the qubit and the drive is:
$H=\omega_{m}a^{\dagger}a+\omega_{q}(t)\frac{\sigma_{z}}{2}+E_{0}(t)\cos[\phi_{d}(t)](\sigma^{-}+\sigma^{+})+g(a^{\dagger}\sigma^{-}+a\sigma^{+}),$ (21)
where we define the instantaneous drive frequency as $\dot{\phi}_{d}(t)=\omega_{d}(t)$. The rapid oscillation on the drive term is removed through a rotating frame transformation by $U=\exp[i\phi_{d}(t)(a^{\dagger}a+\sigma_{z}/2)]$, so that the Hamiltonian becomes (dropping remaining rapidly oscillating terms)
$H_{\mathrm{eff}}=\Delta_{dc}(t)a^{\dagger}a+\frac{\Delta_{dq}(t)}{2}\sigma_{z}+\frac{\Omega(t)}{2}\sigma_{x}+g(a^{\dagger}\sigma^{-}+a\sigma^{+}){\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0},}$ (22)
where $\Delta_{dc}=\omega_{m}-\omega_{d}(t)$ is the cavity-drive detuning, $\Delta_{dq}(t)=\omega_{q}(t)-\omega_{d}(t)$ is the qubit-drive detuning, and $\Omega(t)=E_{0}(t)$. While $\Delta_{dq}(t)$ can be varied by holding the qubit frequency fixed and varying the drive frequency, we observe that the cavity-qubit detuning in this frame, $\Delta_{dc}(t)-\Delta_{dq}(t)=\omega_{m}-\omega_{q}(t)$, can only be varied in time if the physical qubit frequency $\omega_{q}(t)$ is modulated.
@@ Appendix C Protocol Schedule
For both ramp out protocols, we average $10^{4}$ shots to reduce the readout noise from our electronics. The details are discussed in Appendix H.
@@ Appendix D Numerical simulations of the augmented spin-cavity and boosting setups
with $L_{1}=\sqrt{\kappa}a$ (the jump operator modelling photon loss), $L_{2}=\sqrt{\Gamma}\sigma^{-}$ (the jump operator modelling qubit loss, or $T_{1}$ relaxation), and $\{A,B\}=AB+BA$ being the anticommutator. In Sec. 5 we use a similar Lindblad framework in a Hilbert space with two cavities, but with $H_{B}^{+}$ replaced by $H_{\mathrm{boost}}$ and a single jump operator $L_{1}=\sqrt{\kappa_{\mathrm{s}}}a_{\mathrm{s}}$. We solve for $\rho(t)$ given some initial conditions by direct integration in a truncated cavity Hilbert space (we have checked that the state remains far from the truncation throughout the evolution), using the master equation methods of QuTiP Johansson et al. (2013). The data displayed in Figs. 5, 9, and 12 is obtained in this way.
@@ Appendix E Onset of stabilization in fast ramp-out measurements
Figure 10: Onset of stabilization. (a) The probability of being in state |$\ket{e}$ as a function of time for different value of minimum detuning to cavity resonance $\delta/2\pi=(\Delta-B_{0})/2\pi$ measured using the fast ramp out protocol. The orange (blue) curves are obtained by starting with the qubit aligned (anti-aligned) with the field, which we denote as |$\ket{\phi_{+}}_{0}$ and |$\ket{\phi_{-}}_{0}$, respectively. (a.i) Far from resonance ($\delta\gg g$), both state initializations damp out to a mixed state while maintaining a $\pi$-phase shift. (a.ii-a.iii) As the detuning $\delta$ decreases, the state initially prepared in |$\ket{\phi_{+}}$ damps out. Its signal then grows to a non-zero oscillatory amplitude in phase with that from the |$\ket{\phi_{-}}$ state. (a.iv) On resonance, both states are quickly stabilized to |$\ket{\phi_{-}(t)}$. (b) Corresponding data acquired using the slow ramp out protocol, effectively showing the envelope of the oscillatory data in panels (a.i-a.iv). Note that near resonance, the stabilization rate exceeds the oscillation rate leading to significant amplitude variation within a single period, and the slow ramp-out data no longer has a simple correspondence with the envelope of (a). Curves are labeled by the minimum detuning to cavity resonance $\delta/2\pi$ in MHz. Exponential fits to the data are shown in dashed (dotted) lines for preparation in |$\ket{\phi_{-}}$ (|$\ket{\phi_{+}}$). Insets on the right of (a) show a schematic representation of the resonance condition in the $B_{x}-B_{z}$ plane. The solid black line denotes the magnetic field strength and the dashed line denotes the mean qubit-cavity detuning (the colored bands show the ±$\pm g$ width from the cavity-qubit coupling). For both panels, the error bars are smaller than the markers, see appendix H for details.
In addition to the slow ramp-out data presented in Fig. 2, it is also possible to observe the onset of stabilization in the bare qubit |$\ket{g}$, |$\ket{e}$ basis using the fast ramp out protocol.
This protocol gives direct access to the time dependence of the instantaneous eigenstates. States prepared in |$\ket{\phi_{-}}$ (using an initial $\pi$-pulse before the protocol starts) evolve with a $\pi$-phase shift relative to states initialized in |$\ket{\phi_{+}}$ as shown in the blue and orange traces in panels (a.i)-(a.ii) in Fig. 10 respectively. As $\delta=\Delta-B_{0}$ approaches zero, states initially prepared in |$\ket{\phi_{+}}$ damp out before being stabilized in the |$\ket{\phi_{-}}$ state, oscillating at high contrast indefinitely. The instantaneous eigenstate data [shown again in panel (b) for convenience], represents the |$\ket{\phi_{-}}$ population for the oscillatory data presented in panels (a.i)-(a.iv) with 0.5 indicating a perfectly mixed state (i.e. zero oscillation contrast).
@@
Finally, we also observe the stabilization effect for elliptical fields where the resonance condition is only satisfied for a portion of the driving period, as shown in Fig. 11. Here, instead of sweeping the magnitude of $\vec{B}$ to change the detuning, we only sweep the amplitude of the $x$-component of the field, leading to a time-dependent effective detuning between the adiabatic states and the cavity. Just as in the circularly polarized case, when the effective detuning approaches zero, the system is driven to the |$\ket{\phi_{-}}$ state regardless of initial condition and is indefinitely stabilized.
Figure 11: Stabilization with elliptical fields Left: Unnormalized readout amplitude (where a high voltage is proportional to the population in |$\ket{g}$ and a low voltage to that in |$\ket{e}$) as a function of time for different values of the drive field amplitude $B_{x}$, for fixed $B_{z}$ (60 MHz). Blue (orange) traces indicate that the qubit is initialized in |$\ket{\phi_{-}}_{0}$ (|$\ket{\phi_{+}}_{0}$), corresponding to the anti-aligned (aligned) state with respect to the applied field. As the amplitude of $B_{x}$ grows, the quasi-energy states cross resonance with the lossy cavity when $\vec{B}$ points predominantly along $x$, leading to stabilization of the |$\ket{\phi_{-}}$ state. Right: Schematic representation of the applied field (solid black line) and cavity resonance condition (colored dashed line) for different $B_{x}$ strengths (60, 80, 100 MHz) at a fixed cavity detuning $\Delta=90$ MHz.